# Detection of evolutionary shifts in variance under an Ornstein–Uhlenbeck model

### v2

@@ 4.2 Relationship between shift position and detection difficulty
The results indicate that shifts in optima and shifts in variance exhibit distinct detection patterns. For shifts in the optimal value, those occurring near the root or the tips are easier to detect compared to shifts occurring in the middle regions of the tree. In contrast, shifts in variance are generally harder to detect. Our findings show that shifts in variance are more likely to be detected when they occur on branches with a larger number of descendant tips and longer branch lengths, which provides a stronger signal for the variance change. This pattern aligns with our analysis in the Evolutionary shifts in optimal value and variance subsection. The overall results emphasize the importance of the tree structure and branch characteristics in determining the detectability of evolutionary shifts, with shifts in variance requiring more stringent conditions for reliable detection compared to shifts in optima.
@@ 4.3 The influence of estimation of $\alpha$
We conducted simulations to investigate how the misestimation of $\alpha$ affects ShiVa’s shift detection results. In these simulations, we set the true value of $\alpha$ to 1 (corresponding to $\log(\alpha)=0$). To assess the impact of different levels of misestimation, we tested a range of estimated $\alpha$ values, using $\log(\alpha)=-5,-3,-1,0,1,3,5$ as inputs to the ShiVa function. The detection results were evaluated by comparing the predictive log-likelihoods across these different estimated values of $\alpha$.
@@
4.4 The effect of shift position within a branch
@@
Firstly, we conduct a series of simulations with the shift occuring in different positions along a fixed branch. The different locations of the shift are shown in Figure 7 (left). Figure 8 shows the True Positives v.s. False Positives rates and the predictive log-likelihood difference between ShiVa and the true model with shifts in different locations. Location here is a parameter ranging from 0 to 1, with 0 meaning the beginning of the branch and 1 meaning the end of the branch. The ability of the method to detect the shift in variance is not greatly impacted by the position of the shift along the branch. Furthermore, when the shift size is small, the log-likelihood of the model is not much reduced in situations where the location of the shift is not at the beginning of the branch. However, when the magnitude of the shift is larger, the model misspecification does cause a substantial decrease in log-likelihood when the shift is not located at the begining of the branch. Overall, our method is robust to violations of our assumption that shifts occur at the beginning of a branch.
@@
Figure 8: The true positive v.s. false positive rates and the predictive log-likelihood difference of ShiVa with shifts at different locations
Secondly, we simulate scenarios with two shifts on a single branch. If shifts are both positive or both negative, the data will show a stronger signal than when just one shift occurs. Therefore, we simulate situations where two opposite shifts occur (at locations = 0 and 1) (Figure 7 right). To compare, we also simulate situations where only one shift occurs at the beginning of that branch. Table 2 shows that when the two opposite shifts occur on that branch, our method cannot accurately detect the shift in variance on that branch and the false positive rate of shifts in optimal value increases. In this case, the violation of the assumption causes some difficulties for our method. However, in the general case, it is not a major concern.
@@
5 Case study: Cordylid data
Cordylidae is a family of lizards in Sub-Sarharan Africa (Stanley, Bauer, Jackman, Branch, and Mouton, 2011). Broeckhoven et al. (2016a) selected 28 species of Cordylidae, including five outgroup taxa, and reconstructed the phylogenetic tree based on their DNA sequences using BEAST. They also collected 20 morphological traits, including spine, limb and head dimensions and used pPCA (Revell, 2009) to perform dimension reduction. Based on their analysis, the first principal component pPCA1 represented a gradient from highly armored species to lightly armored species. The phylogenetic tree and principal component data are available in the R package phytools: http://www.phytools.org/Rbook/. We used the cordylid tree and the first principal component they provided to compare the shift detection results of $\ell$1ou, phyloEM and ShiVa.
Figure 9 shows the results given by different methods. ShiVa uses cross validation to tune the hyperparameter $\lambda_{1}$, therefore the final results have randomness. We repeat the method 3 times and select the best model based on BIC for this real data analysis. The model estimated by ShiVa gives the highest log-likelihood (18.2) and the lowest BIC (40.24). $\ell$1ou+pBIC and phyloEM detected limited numbers of shifts, with lower log-likelihood and higher BIC compared to the other two methods. Comparing the results of $\ell$1ou+BIC and ShiVa, $\ell$1ou+BIC detected branches 28, 43, 44 as shifts in optimal value, and ShiVa detected a shift in variance: branch 47, without detecting the above shifts in optimal value. From the results of ShiVa, it is possible that the subgroup Cordylus has a different variance from other taxa. This interesting finding may be neglected, and indeed detected as multiple false shifts in optimal values using the other methods.
To determine which of the shifts in optimal value and variance selected by the various methods are more plausible, we simulated 100 datasets based on each selected model and applied $\ell$1ou, phyloEM and ShiVa to check whether the models selected by the different methods are plausible for this combination of true shifts. Table 3 shows the validation results. If the models selected by $\ell$1ou+pBIC or by phyloEM were true, then it is very unlikely that $\ell$1ou+BIC or ShiVa would detect as many shifts as they do for the real data. Similarly, if the model selected by $\ell$1ou+BIC were true, it would be unlikely that ShiVa would select so many shifts in variance. On the other hand, the results for the simulation based on the shifts identified by ShiVa are more similar to the real data results. Figure 11 shows more detailed results from these simulations. We see that in terms of the total number of shifts selected in both optimal value and variance, the simulation results for the model selected by ShiVa are far more similar to the real data. This suggests that this model is more likely to be close to the truth. We see that ShiVa often selects surrogates to the true shifts in simulations, so we cannot be confident of the exact location of the shifts, but it certainly appears that there are shifts in variance, and these shifts have impacted the performance of shift detection methods that only detect shifts in optimal value. The shifts in optimal value detected by these methods, but not by ShiVa are likely to be false positives caused by the shifts in variance.
Considering the log-likelihood, BIC, and the validation results, the shifts in optimal value and variance detected by ShiVa in the cordylid dataset fit the data better compared to other methods.
Figure 9: The shift detection results of cordylid data of $\ell$1ou+pBIC (left top), $\ell$1ou+BIC (right top), phyloEM (left bottom) and ShiVa (right bottom)
Table 3: True positive versus False positive rates of different methods using data simulated from the estimated model given by Figure 9
Model simulated from Shifts in optimal value Shifts in variance Method True positive (optimal value) False positive (optimal value) True positive (variance) False positive (variance) l1ou_pBIC 17;50 ShiVa 2.00 0.24 0.00 0.63 l1ou+pBIC 2.00 0.05 0.00 0.00 l1ou+BIC 2.00 4.80 0.00 0.00 phyloEM 2.00 0.13 0.00 0.00 l1ou_BIC 3;4;10;17;25;28;43;44;50 ShiVa 3.43 0.87 0.00 0.14 l1ou+pBIC 2.32 0.01 0.00 0.00 l1ou+BIC 7.18 3.67 0.00 0.00 phyloEM 3.53 1.19 0.00 0.00 phyloEM 3;17;50 ShiVa 2.62 1.16 0.00 0.71 l1ou+pBIC 2.39 0.05 0.00 0.00 l1ou+BIC 2.87 5.62 0.00 0.00 phyloEM 2.65 0.20 0.00 0.00 ShiVa 3;10;17;18;26;49;50 16;17;47 ShiVa 5.94 3.08 0.40 0.82 l1ou+pBIC 1.50 0.02 0.00 0.00 l1ou+BIC 5.72 6.56 0.00 0.00 phyloEM 3.48 0.78 0.00 0.00
Figure 10: Left: True positive versus false positive rates of the shifts in optimal values detected by different methods. Right: False positive rates of the shifts in variance detected by ShiVa.
Figure 11: The selecting frequencies of shifts given by different methods; deeper color shows higher selecting frequency. (Green: selecting frequencies of shifts in optimal value; Red: selecting frequencies of shift in variance). The row names show the model data are simulated from; the column names show the method applied to the generated data. The asterisk marks show the true shifts of the model that the data are simulated from.
6 Conclusion
Our results showed that ShiVa effectively balances True Positives (TP) and False Positives (FP) across different scenarios. When there was only a shift in the optimal value, ShiVa performed comparably to other methods, such as $\ell$1ou and phyloEM, maintaining similar predictive accuracy while achieving a better balance between TP and FP compared to PCMFit, which tended to overfit in weak signal scenarios. When a shift in variance was present, ShiVa demonstrated a significant advantage over other methods, achieving higher predictive log-likelihood and effectively distinguishing between variance and mean shifts. This led to fewer False Positives compared to $\ell$1ou and phyloEM, which often misinterpreted variance shifts as multiple shifts in optimal value. Although PCMFit achieved higher TP rates, it generated significantly more FP, indicating overfitting.
We applied our method to the first principal component of trait values for the cordylid data from Broeckhoven et al. (2016a). ShiVa detected 7 shifts in optimal value and 3 shifts in variance and the model given by ShiVa has the highest log-likelihood and lowest BIC compared to $\ell$1ou and phyloEM. The validation result also shows that Shiva would be unlikely to estimate the model it did if any of the models selected by other methods were true.
For model simplicity, we assume that the variance parameter is constant along each branch and that all shifts in variance happen at the beginning of the branch. Simulations of cases where these assumptions are violated show that our method is robust to misspecification in these assumptions. An interesting future research direction is to extend our model to allow shifts in variance at internal positions along a branch. This would allow us to estimate the exact time of a shift in variance. Writing the likelihood for this case is straightforward, but more work may be needed to ensure stability of the estimates. Another direction for future work is to improve the computational efficiency of the method. The likelihood calculation involves a large number of matrix inverse computations,which can be computationally expensive for large trees. Bastide et al. (2021) provide an efficient algorithm to calculate the log likelihood and its derivatives for certain phylogenetic models. If this method could be adapted to our method, it would lead to a major improvement in the computational speed, allowing our method to scale to larger phylogenies.
Appendix A Supplemental Materials
f
@@ Appendix A Supplemental Materials
Average Computation Time Shift Position (mean) Beta Shift Position (variance) Gamma ShiVa l1ou-pBIC l1ou-BIC phyloEM PCMFit 43,71,192 1 0 0 86.80 7.66 8.43 69.64 11561.58 43,71,192 3 0 0 74.95 8.37 8.36 68.09 27477.04 43,71,192 5 0 0 88.87 8.76 8.86 68.05 36900.89 43,71,192 -1 0 0 71.88 7.99 8.40 69.21 13952.25 43,71,192 -3 0 0 67.89 8.61 8.56 70.58 35867.14 43,71,192 -5 0 0 76.54 9.13 9.16 72.69 40202.81 6,71,197,3,88,191,98 1 0 0 117.94 7.80 8.05 70.53 14863.06 6,71,197,3,88,191,98 3 0 0 80.94 6.06 6.08 70.63 31135.83 6,71,197,3,88,191,98 5 0 0 81.68 8.55 8.86 67.94 59322.66 6,71,197,3,88,191,98 -1 0 0 83.20 7.79 8.04 69.02 15652.61 6,71,197,3,88,191,98 -3 0 0 98.53 7.42 6.84 69.47 29763.69 6,71,197,3,88,191,98 -5 0 0 69.55 7.99 7.85 70.15 60076.49 0 0 196 1 68.06 8.08 8.35 69.46 14515.10 0 0 196 1.5 72.20 8.03 8.31 69.51 15649.65 0 0 196 2 83.65 8.39 8.54 70.18 17266.36 0 0 196 3 90.27 7.61 8.14 69.24 18907.37 0 0 196 5 89.08 8.27 8.42 69.61 18364.71 0 0 196 7 101.44 8.03 8.51 67.09 18858.03 0 0 196 -1 99.09 8.26 8.66 67.89 16773.86 0 0 196 -1.5 113.54 8.21 8.72 67.30 19419.61 0 0 195,130 1,-0.5 80.75 8.37 8.68 70.94 16275.18 0 0 195,130 3,-1 93.39 8.11 8.39 70.21 18297.14 0 0 195,130 5,-1.5 159.13 12.06 12.58 116.55 36602.57 0 0 195,130 7,-1.5 114.15 8.60 8.86 69.08 22939.48 0 0 195,130 1 85.99 8.15 8.31 70.17 15881.60 0 0 195,130 3 91.66 8.10 8.38 70.03 17969.21 0 0 195,130 5 104.25 8.69 9.17 71.10 25327.97 0 0 195,130 7 95.08 8.63 8.94 69.17 23545.64
Appendix B Data accessibility
The tree of Anolis lizards is provided by Mahler et al. (2013b). It can be found in the Dryad data repository: https://doi.org/10.5061/dryad.9g182. The phylogenetic tree and principal component data of cordylid data (Broeckhoven, Diedericks, Hui, Makhubo, and Mouton, 2016b) can be found in the Dryad data repository: https://doi.org/10.5061/dryad.k186f. Our R package is available at https://github.com/WenshaZ/ShiVa.
Appendix C Author contributions
Appendix D Acknowledgement
LSTH was supported by the Canada Research Chairs program, the NSERC Discovery Grant RGPIN-2018-05447, and the NSERC Discovery Launch Supplement. TK was supported by the NSERC Discovery Grant RGPIN/4945-2014.

### v3

@@ 4.2 Relationship between shift position and detection difficulty
The results indicate that shifts in optima and shifts in variance exhibit distinct detection patterns. Shifts in the optimal value tend to be more detectable when they occur near the root or at the tips, while those occurring in the middle regions of the tree are generally harder to identify. In contrast, shifts in variance are more challenging to detect overall. Our findings show that shifts in variance are more likely to be detected when they occur on branches with a greater number of descendant tips and longer branch lengths, which amplify the signal of the variance change. Notably, the two branches immediately below the root node consistently show lower detection probabilities for variance shifts. This is not due to a lack of signal, but because shifts on these two branches are statistically non-identifiable. This observation is consistent with our analysis in the Evolutionary shifts in optimal value and variance subsection. Overall, these results highlight how the structure of the phylogeny and the position of shifts critically influence the detectability of evolutionary changes, with variance shifts requiring more informative configurations for reliable inference compared to shifts in optima.
@@ 4.3 The influence of estimation of $\alpha$
We conducted simulations to examine how misestimating $\alpha$ affects ShiVa’s ability to detect shifts. The true value of $\alpha$ was set to 1 (i.e., $\log(\alpha)=0$). To assess the impact of varying estimates, we ran ShiVa using a range of fixed $\alpha$ values corresponding to $\log(\alpha)=-5,-3,-1,0,1,3,5$. We evaluated the results by comparing predictive log-likelihoods across these settings. Importantly, the predictive log-likelihood was computed using the true $\alpha$ value to ensure fair comparison of the detected shifts and to avoid confounding effects from the misestimated $\alpha$.
@@
5 Case study
To illustrate the application of our method, we analyzed two empirical datasets. The first dataset concerns floral diameter in the parasitic plant family Rafflesiaceae, using the phylogeny flowerTree from Davis et al., (2007). The second dataset involves buccal morphology in Centrarchidae sunfishes, based on a time-calibrated phylogeny of 28 species from Revell and Collar, (2009). We applied ShiVa to log-transformed floral diameter in the Euphorbiaceae dataset and log-transformed buccal length in the Centrarchidae dataset, comparing results with $\ell$1ou+pBIC, $\ell$1ou+BIC, PhylogeneticEM, and PCMFit. Notably, ShiVa incorporates cross-validation during parameter tuning, introducing some randomness into the results. To ensure robustness, we recommend running ShiVa three times and selecting the model with the lowest BIC, as we did in our empirical analyses.
Figure 8 presents the shift detection results on the flower dataset. ShiVa identified one shift in the optimal value and one shift in evolutionary variance, with estimated values of $\sigma^{2}=0.33$ and $\gamma=2.42$, respectively. Compared to the results of $\ell$1ou+pBIC, ShiVa detects one additional variance shift. The shift size of 2.42 is substantial, especially when considered relative to the original variance, highlighting a significant change in evolutionary dynamics.
Although ShiVa appears to yield a lower log-likelihood and higher BIC than $\ell$1ou+pBIC, this comparison includes the estimation of the selection strength parameter $\alpha$. When comparing detection results under a fixed $\alpha$, ShiVa achieves a higher log-likelihood and lower BIC than $\ell$1ou+pBIC, suggesting that modeling variance shifts improves model fit. Notably, compared to $\ell$1ou+pBIC, ShiVa detects an additional variance shift on branch 16. To evaluate the significance of this shift, we performed a likelihood ratio test. The log-likelihood of the model with only a mean shift on branch 41 is $-21.8940$, while the model that also includes a variance shift on branch 16 achieves a log-likelihood of $-18.3639$. The resulting test statistic corresponds to a $p$-value of $0.0293$, which is below the $0.05$ threshold, indicating that the variance shift on branch 16 is statistically significant. PhyloEM and $\ell$1ou+BIC yield identical shift detections and attain the highest log-likelihood among all methods.
To better understand the behavior of each method, we grouped the detected models into four representative types: (1) ShiVa: one shift in optimal value and one shift in variance; (2) $\ell$1ou+pBIC: one shift in optimal value; (3) $\ell$1ou+BIC / PhyloEM: three shifts in optimal value; (4) PCMFit: changes in all parameters on two shifted branches. To assess how well each method performs when a given model is the true generative process, we simulate 100 datasets from each estimated model and re-apply all five methods to the simulated data. We then compute the predictive log-likelihood to compare model fit across methods. As shown in Table 2, ShiVa consistently achieves relatively high and stable predictive log-likelihood, even when the data are generated from models estimated by other methods. This suggests that ShiVa produces robust and reasonable detection results across a variety of underlying evolutionary scenarios, supporting its credibility and generalizability.
Figure 8: The detection results on the flower dataset using different methods are shown. Green circles indicate shifts in optimal trait values, while pink squares denote shifts in evolutionary variance. The numbers inside the markers correspond to the node indices where shifts occur, and the colored numbers next to them represent the estimated magnitudes of those shifts.
Table 2: The median predictive log-likelihood of different methods, evaluated on datasets simulated from each of the estimated models. (flower dataset)
Model Simulated ShiVa $\ell$1ou+pBIC $\ell$1ou+BIC phyloEM PCMFit True
(1) -23.8826 -28.8382 -34.8087 -28.6572 -167.6036 -18.3726
(2) -21.2532 -23.0154 -29.6700 -23.0154 -153.1015 -18.0685
(3) -20.0300 -21.1384 -18.1560 -20.6989 -234.6207 -7.0669
(4) -23.2938 -26.3329 -70.2521 -26.1872 -284.8035 -13.4989
Figure 9 summarizes the results from the In the Centrarchidae analysis, ShiVa detected two shifts in the optimal trait value and three shifts in evolutionary variance. It estimated a saturated model by assigning an extremely small value to the original $\sigma^{2}$. Among the three variance shifts, the one on branch 48—with a magnitude of 0.309—is the most substantial, while the other two primarily restore typical levels of diffusion variance. $\ell$1ou+pBIC identified two shifts in the optimal trait value, while $\ell$1ou+BIC inferred 11 such shifts. PhylogeneticEM and PCMFit detected no shifts under their respective model selection criteria.
ShiVa achieved a higher log-likelihood than $\ell$1ou+pBIC, PhylogeneticEM, and PCMFit. Although $\ell$1ou+BIC attained the highest log-likelihood, it did so by detecting a large number of shifts in the optimal value—including several within the clade descending from the branch where ShiVa detected a variance shift. For example, ShiVa detected a variance shift on branch 48 with a substantial magnitude of 0.309. Under that branch, $\ell$1ou+BIC detected two shifts in the optimal value (branches 29 and 31), and $\ell$1ou+pBIC also detected a shift at branch 29. Despite its higher log-likelihood, $\ell$1ou+BIC resulted in a higher BIC than ShiVa.
Using this dataset, we grouped the estimated models into four representative types: (1) ShiVa: 2 shifts in optimal value and 3 shifts in variance; (2) $\ell$1ou+pBIC: 2 shifts in optimal value; (3) $\ell$1ou+BIC: 11 shifts in optimal value; (4) PhyloEM / PCMFit: no shifts detected.
As in the previous analysis, we simulated 100 datasets from each estimated model and re-applied all five methods to the simulated data. We then computed the predictive log-likelihood to evaluate model fit across methods. As shown in Table 3, ShiVa consistently achieves relatively high and stable predictive log-likelihoods, except for model (2). Its performance is affected in scenarios with a large number of shifts in the optimal value.
Figure 9: The detection results on the centrarchidae data using different methods are shown. Green circles indicate shifts in optimal values, and pink squares indicate shifts in evolutionary variance. The numbers inside the markers represent the branch indices where shifts occur, while the colored numbers next to them denote the estimated shift magnitudes.
Table 3: The median predictive log-likelihood of each method, evaluated on simulated datasets generated from each estimated model. (sunfish dataset)
Model Simulated ShiVa $\ell$1ou+pBIC $\ell$1ou+BIC phyloEM PCMFit True
(1) 27.7391 17.7871 2.2102 19.3929 19.9311 72.3057
(2) 43.2906 49.1707 43.7983 41.2296 40.9689 57.3213
(3) 25.9643 49.8254 85.6678 39.0104 46.2593 99.8206
(4) 41.8720 41.3245 31.9973 43.1736 44.7510 46.0488
6 Conclusion
Our results showed that ShiVa effectively balances True Positives (TP) and False Positives (FP) across different scenarios. When there was only a shift in the optimal value, ShiVa performed comparably to other methods, such as $\ell$1ou and phyloEM, maintaining similar predictive accuracy while achieving a better balance between TP and FP compared to PCMFit, which tended to overfit in weak signal scenarios. When a shift in variance was present, ShiVa demonstrated a significant advantage over other methods, achieving higher predictive log-likelihood and effectively distinguishing between variance and mean shifts. This led to fewer False Positives compared to $\ell$1ou and phyloEM, which often misinterpreted variance shifts as multiple shifts in optimal value. Although PCMFit achieved higher TP, it generated significantly more FP, indicating overfitting.
Our empirical analyses on floral diameter in Euphorbiaceae and buccal morphology in Centrarchidae sunfishes demonstrate the practical utility of ShiVa. In both datasets, ShiVa successfully identified meaningful variance shifts that were missed or misinterpreted by methods focusing only on optimal value shifts. While ShiVa may yield slightly lower likelihoods in some cases, it achieves better model fit when accounting for variance shifts and maintains strong predictive performance across models. These results underscore ShiVa’s robustness and its ability to capture complex evolutionary dynamics that may be overlooked by existing approaches.
For model simplicity, we assume that the variance parameter is constant along each branch and that all shifts in variance happen at the beginning of the branch. Simulations of cases where these assumptions are violated show that our method is robust to misspecification in these assumptions. An interesting future research direction is to extend our model to allow shifts in variance at internal positions along a branch. This would allow us to estimate the exact time of a shift in variance. Writing the likelihood for this case is straightforward, but more work may be needed to ensure stability of the estimates. Another direction for future work is to improve the computational efficiency of the method. The likelihood calculation involves a large number of matrix inverse computations,which can be computationally expensive for large trees. Bastide et al., (2021) provide an efficient algorithm to calculate the log likelihood and its derivatives for certain phylogenetic models. If this method could be adapted to our method, it would lead to a major improvement in the computational speed, allowing our method to scale to larger phylogenies.
7 Data accessibility
The phylogenetic tree of Anolis lizards is provided by Mahler et al., 2013a and it can be accessed via the R package l1ou (https://github.com/khabbazian/l1ou). The phylogenetic tree and trait data for the Euphorbiaceae flower dataset (Davis et al.,, 2007) are available through the R package phylolm, and the Centrarchidae sunfish dataset (Revell and Collar,, 2009) can be accessed via the R package phytools. Both packages are available on CRAN. Our R package ShiVa is available at https://github.com/WenshaZ/ShiVa, and the simulation and case study code can be found at https://github.com/WenshaZ/ShiVa-Experiments.
8 Author contributions
9 Acknowledgement
LSTH was supported by the Canada Research Chairs program, the NSERC Discovery Grant RGPIN-2018-05447, and the NSERC Discovery Launch Supplement. TK was supported by the NSERC Discovery Grant RGPIN-2023-03332.
Supplemental Materials
A The effect of shift position within a branch
@@
Firstly, we conduct a series of simulations with the shift occuring in different positions along a fixed branch. The different locations of the shift are shown in Figure 10 (left). Figure 11 shows the True Positives v.s. False Positives and the difference of predictive log-likelihood between ShiVa and the true model with shifts in different locations. Location here is a parameter ranging from 0 to 1, with 0 meaning the beginning of the branch and 1 meaning the end of the branch. The ability of the method to detect the shift in variance is not greatly impacted by the position of the shift along the branch. Furthermore, when the shift size is small, the log-likelihood of the model is not much reduced in situations where the location of the shift is not at the beginning of the branch. However, when the magnitude of the shift is larger, the model misspecification does cause a substantial decrease in log-likelihood when the shift is not located at the begining of the branch. Overall, our method is robust to violations of our assumption that shifts occur at the beginning of a branch.
@@
Figure 11: The true positive v.s. false positive and the predictive log-likelihood difference of ShiVa with shifts at different locations
Secondly, we simulate scenarios with two shifts on a single branch. If shifts are both positive or both negative, the data will show a stronger signal than when just one shift occurs. Therefore, we simulate situations where two opposite shifts occur (at locations = 0 and 1) (Figure 10 right). To compare, we also simulate situations where only one shift occurs at the beginning of that branch. Table 4 shows that when the two opposite shifts occur on that branch, our method cannot accurately detect the shift in variance on that branch and the false positive number of shifts in optimal value increases. In this case, the violation of the assumption causes some difficulties for our method. However, in the general case, it is not a major concern.
@@
B Supplementary figures and tables
@@
Average Computation Time (seconds) Shift Position (mean) Beta Shift Position (variance) Gamma ShiVa l1ou-pBIC l1ou-BIC phyloEM PCMFit 43,71,192 1 0 0 86.80 7.66 8.43 69.64 11561.58 43,71,192 3 0 0 74.95 8.37 8.36 68.09 27477.04 43,71,192 5 0 0 88.87 8.76 8.86 68.05 36900.89 43,71,192 -1 0 0 71.88 7.99 8.40 69.21 13952.25 43,71,192 -3 0 0 67.89 8.61 8.56 70.58 35867.14 43,71,192 -5 0 0 76.54 9.13 9.16 72.69 40202.81 6,71,197,3,88,191,98 1 0 0 117.94 7.80 8.05 70.53 14863.06 6,71,197,3,88,191,98 3 0 0 80.94 6.06 6.08 70.63 31135.83 6,71,197,3,88,191,98 5 0 0 81.68 8.55 8.86 67.94 59322.66 6,71,197,3,88,191,98 -1 0 0 83.20 7.79 8.04 69.02 15652.61 6,71,197,3,88,191,98 -3 0 0 98.53 7.42 6.84 69.47 29763.69 6,71,197,3,88,191,98 -5 0 0 69.55 7.99 7.85 70.15 60076.49 0 0 196 1 68.06 8.08 8.35 69.46 14515.10 0 0 196 1.5 72.20 8.03 8.31 69.51 15649.65 0 0 196 2 83.65 8.39 8.54 70.18 17266.36 0 0 196 3 90.27 7.61 8.14 69.24 18907.37 0 0 196 5 89.08 8.27 8.42 69.61 18364.71 0 0 196 7 101.44 8.03 8.51 67.09 18858.03 0 0 196 -1 99.09 8.26 8.66 67.89 16773.86 0 0 196 -1.5 113.54 8.21 8.72 67.30 19419.61 0 0 195,130 1,-0.5 80.75 8.37 8.68 70.94 16275.18 0 0 195,130 3,-1 93.39 8.11 8.39 70.21 18297.14 0 0 195,130 5,-1.5 159.13 12.06 12.58 116.55 36602.57 0 0 195,130 7,-1.5 114.15 8.60 8.86 69.08 22939.48 0 0 195,130 1 85.99 8.15 8.31 70.17 15881.60 0 0 195,130 3 91.66 8.10 8.38 70.03 17969.21 0 0 195,130 5 104.25 8.69 9.17 71.10 25327.97 0 0 195,130 7 95.08 8.63 8.94 69.17 23545.64
(a) $\alpha$ = 0.01
(b) $\alpha$ = 0.1
(c) $\alpha$ = 10
Figure 14: Comparison of True positives and False positives of different methods (Varying $\alpha$)